A uniform electric field of 10 N/C exists in the vertically downward direction. Find the increase in the electric potential as one goes up through a height of 50 cm.

Grade:10

2 Answers

Jitender Pal
9 years ago
Sol. . E = 10 n/c, S = 50 cm = 0.1 m E = dV/dr or, V = E * r = 10 * 0.5 * 5 cm
Kevin Nash
9 years ago
Sol. E = 10 n/c, S = 50 cm = 0.1 m E = dV/dr or, V = E * r = 10 * 0.5 * 5 cm